Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
fiori, paoletti, tringali, washimi, sugimoto, karathanasis, romero, menendez - 18:51 Saturday 05 October 2019 (47107) Print this report
Task ENV02: investigation of web 40-45Hz peaks

Commissioning task ENV02. PURPOSE: investigate the bunch of Hrec peaks between 41 and 43Hz (more precisely 41.15, 41.9, 42.4, 42.9 Hz). See first attached plot. Previous investigations in 46037 (peaks have some small coherence with WEB magnetometers and UPS).

We sniffed with portable magnetometer (tmp channel ENV_WEB_MAG) various locations (PCAL rack, PCAL boxes placed on tower base at East side, LC SE corner, camera NE corner) but found no evidence of magnetic emissions associated to the peaks.

We tapped various components, see attached log file with times.

  • We saw no effect in DARM when tapping: SWEB link pipe, SWEB vacuum chamber, HWFS bench pipe, Cryotrap bellows, Cryotrap tank (we hammered on the joints of lateral rods with the help of Antonio Pasqualetti).
  • We saw an effect when tapping on "costolone": some bumps apeared in Hrec in correspondence with loud vibration peaks seen by the vacuum chamber ENV accelerometer, see Figure #2.  But nothing at 41-43Hz. The frequency pattern resembles scattered light. The coupling seems not critical, but worth to ty to extract estimated coupling from the data.

TOWER VACUUM CHAMBER EXCITATIONS, using Kagra portable shaker
We attached one accelerometer to WE vacuum chamber at the mid stage (second ladder): ENV_WE_TOP_ACC_X (same used for 46534). Then:

  • *Without ITF* we knocked on the virola and clearly excited resonances at 41.35 and 70.3Hz: Figure #3 and Figure #4. We repeated using the Kagra bluetooth shaker attached to the tower mid stage virola: Figure #5
  • *With interferometer* we injected some noise using kagra's shaker this time attached at the bottom flange (W):
    • 41.3Hz line (to match the chamber mode): Hrec excites easily
    • recursive sweep between 35Hz and 45Hz: we see Hrec excites each time the sweep frequency crosses 41.3-41.4Hz but not elsewere: Figure #6
    • 70.3 hz line (to match the other chamber mode): Hrec excites easily
    • recursive sweep between 65Hz and 75Hz: we see Hrec excites each time the sweep freq. matches exactly 70.3Hz: Figure #7


Since we noticed the 41-43Hz peaks have some (small) coherence with PCAL_WEB_PSD2 signals (but no peaks in this channels), we want to try to have it off. (note the WEB PCAL was disabled since Sept.5 because of a failure). We contacted Dimitri Esteves (PCAL on call) who first set to zero the voltage sent to the Pcal PSD and then advised to disconnect the the 4 DAC cables of the WE PCAL PSD1 and PSD2 driving, which we did. We lesft the cables disconnected. Yet the 41-43 Hz peaks today (Oct 5) are still there.


  • the source of the 41-43 Hz is still not found
  • we understood the origin of the peaks that we excited when injecting acoustic noise at WEB (see 46708 ): they are vibration modes of the vacuum chamber. at least the 70.30Hz peak is present in the quiet sensitivity (see so it is worth to further investigate the coupling. Couls it be scattered light? However the chamber mode at 41.35 Hz seems not to be one of the 41-43Hz family, but it has to be better evaluated.
  • we find some scattered light (tipical bumps pattern) when knocking on the costolone, which seems not critical, but should be better quantified.

TO DO: the magnetic sniffing of 41-43Hz source can continue also without ITF, i.e. next TUESDAY morning.


Images attached to this report
Non-image files attached to this report
Comments to this report:
washimi - 20:02 Tuesday 08 October 2019 (47155) Print this report

I check the current status of WEB 4-peaks in 40-45Hz. (Please see the plots.)

What I found and my guess:

  • Hrec is almost same to  Sep.30 (43.0Hz peak moved to 42.86Hz)
  • The coherence between MAG & UPS is not peak, but bump.
    • It is nancence to search these magnetic-field lines.
    • One possible story is the magnetic-field excites something modes of suspension
  • Though the coherence between PCAL & Hrec / PCAL & MAG becomes higher (factor 2-3), Hrec is not change bery much. 
    • PCAL is not the origin of these line noise.


Therefore, I recommend to do a magnetic injection in detail. (White spectrum or swept sine)

Images attached to this comment
washimi - 10:57 Thursday 10 October 2019 (47177) Print this report

more investigation :

(1) FFTs for all DOF in UPS, IPS, and MAG  --> UPS-S, MAG-N, MAG-W, MAG-V have the bump.

(2) Wider lange FFTs and COHE for UPS-S and MAG-N --> Also 55-63Hz, 35-85Hz, 90-95Hz, and 98-99.5Hz

Images attached to this comment
Paoletti - 12:34 Thursday 10 October 2019 (47179) Print this report

In the attached plot I show a comparison of some WEB sensors before (purple) and after (blue) that the auxiliary UPS (faulty in Aug. 27, due to a mouse) was put back in operation (September the 3rd).

It seems that something has changed on the "S" line: there are very wide sidebands that pop up (+/-7Hz from the 50Hz mains) not present before. These sidebands create a magnetic field that is sensed by all three magnetic sensors that are very close to the base of the tower.

I've also checked the old data, it seems this noise was not present before Sep. the 3rd.

Hard to say if it is a noise generated by the second UPS, or if it is due to a load on the "S" line. We should do a selective shutdown of both UPSs (once at a time) just to better understand what is happening.


Images attached to this comment
Paoletti - 16:39 Monday 14 October 2019 (47217) Print this report

Today we have switched off for a few minutes the second UPS (the one repaired and put back into operation from 3 September); all the magnetic noise around the 50Hz power line (and more!) has disappeared.

We left it on again for reliability reasons, the company (Riello) has been informed, a technician should come next week to have a look.

(UTC timing of the switch off: good data from 13:03:30 to 13:09:30)

Images attached to this comment
Search Help